OpenStudy (anonymous):

how do i subtract this??? 81degrees24'15" 42degrees30'10"

OpenStudy (anonymous):

start by looking at the seconds 15-10=?? then the minutes 24-30= realize that this will be a negative number so subtract a degree from 81 to get (24+60)-30=? now look at the degrees 80-42=?

OpenStudy (anonymous):

so it makes sense?

OpenStudy (anonymous):

yes, thank you very much!!